OpenStudy (wintersomerville22):

The lengths of the legs of a right triangle are 8 inches and twelve inches. What is the exact length of the hypotenuse? 4 inches the square root of twenty inches the square root of forty inches the square root of two hundred eight inches

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge a^2 + b^2 = c^2$$

OpenStudy (skullpatrol):

Let a = 8 and b = 12, then solve for c.

OpenStudy (skullpatrol):

$$\huge 8^2 + 12^2 = c^2$$

OpenStudy (skullpatrol):

$$64 + 144 = c^2$$